# BALLAST: Bayesian Active Learning with Look-ahead Amendment for Sea-drifter Trajectories under Spatio-Temporal Vector Fields

### v1

@@ Abstract
We introduce a formal active learning methodology for guiding the placement of Lagrangian observers to infer time-dependent vector fields – a key task in oceanography, marine science, and ocean engineering – using a physics-informed spatio-temporal Gaussian process surrogate model. The majority of existing placement campaigns either follow standard ‘space-filling’ designs or relatively ad-hoc expert opinions. A key challenge to applying principled active learning in this setting is that Lagrangian observers are continuously advected through the vector field, so they make measurements at different locations and times. It is, therefore, important to consider the likely future trajectories of placed observers to account for the utility of candidate placement locations. To this end, we present BALLAST: Bayesian Active Learning with Look-ahead Amendment for Sea-drifter Trajectories. We observe noticeable benefits of BALLAST-aided sequential observer placement strategies on both synthetic and high-fidelity ocean current models.
@@ 1 Introduction
We propose BALLAST — Bayesian Active Learning with Lookahead Amendment for Sea-drifter Trajectories — a methodology designed to place Lagrangian observers. BALLAST accounts for the data structure of Lagrangian observers by simulating hypothetical trajectories using vector fields. Here, we used a spatio-temporal vector-output Gaussian process (GP) surrogate model (see Figure 1 for an illustration) and an information-theoretic utility function for the active learning, and devised an original application of the stochastic partial differential equation (SPDE) approach of GP (Sarkka et al., 2013) for efficient implementations of BALLAST.
Our contributions can be summarised as follows: (i) we introduce active learning concepts to the literature of Lagrangian observer placements, (ii) we propose BALLAST, a novel active learning amendment that accounts for Lagrangian observations using samples from surrogates, and (iii) we develop a new GP posterior sampling method combining standard GP regression and the SPDE approach for efficient BALLAST utility computation, which may be of independent interest especially for scenarios using spatio-temporal GP models with non-gridded observations. Our numerical results suggest noticeable benefits of BALLAST-aided active learning for sequential observer deployment on both synthetic and high-fidelity ocean current models.
@@ 2.1 Sequential Experimental Design
Sequential experimental design, with active learning being a special case, selects an optimal measurement point $x^{*}_{n}$ from measurement set $X$ at each time $t_{n}$ using existing data $\mathcal{D}_{n}$ by optimising the expectation of utility function $U$ over the posterior predictive distribution $p(y|\mathcal{D}_{n},x)$ at $x\in X$, mathematically formulated as
@@
In our spatio-temporal setting, the object of interest $f$ depends on both space and time. We assume the measurement times are predetermined, so we only select the placement location of the Lagrangian observers at each measurement time. This assumption simplifies the sequential experimental design search space while maintaining sufficient realisms of real-world drifter deployment practices (Lilly and Pérez-Brunius, 2021).
2.2 Gaussian Process for Time-Dependent Vector Fields
While the literature focuses primarily on a static but unknown object of interest $f$, commonly modelled by a GP (Williams and Rasmussen, 2006), in this study the system of interest is a time-dependent vector field. We therefore consider a spatio-temporal, vector-output GP as the surrogate used for active learning.
Following Ponte et al. (2024), we consider an extension to the Helmholtz kernel $k_{\text{Helm}}$ of Berlinghieri et al. (2023) – a vector-output kernel (Alvarez et al., 2012) using the Helmholtz decomposition (Bhatia et al., 2012) to more realistically portray the vector field structure – by including a separable temporal kernel $k_{\text{time}}$, which results in the temporal Helmholtz kernel $k_{\text{tHelm}}\left((\bm{s},t),(\bm{s}^{\prime},t^{\prime})\right)=k_{\text{Helm}}(\bm{s},\bm{s}^{\prime})k_{\text{time}}(t,t^{\prime})$ with $\bm{s},\bm{s}^{\prime}\in\operatorname{\mathbb{R}}^{2},t,t^{\prime}\in\operatorname{\mathbb{R}}$. The spatial Helmholtz kernel $k_{\text{Helm}}$ is constructed as a linear combination of the potential kernel and the stream function kernel, which are two two-dimensional input, scalar output kernels such as the SE kernel (see Section B.2 for details). The temporal kernel $k_{\text{time}}$ is set to be a Matérn $3/2$ kernel: evidence in oceanographic research suggests that the smoothness $\nu$ is around 2; as this leads to a non-analytic representation of the Bessel function, often $\nu=3/2$ is taken as a sufficient approximation (Lilly et al., 2017; Ponte et al., 2024). We will also use $\bm{x}=(\bm{s},t)$ to denote the input of the GP. An illustration of the spatio-temporal GP regression of Lagrangian trajectories is displayed as Figure 1.
@@ 3 Active Learning of Time-Dependent Vector Fields
and $\sigma_{\text{obs}}$ is the standard deviation of the observation noise. The active learning selects the placement location of Lagrangian observers at each placement time in order to maximise the information gain of the full vector field surrogate $f$.
@@
Let $t_{n}$ be the time when we are deciding the placement location of the $n$-th observer. The observations so far are denoted by $\mathcal{D}_{n}=\{X_{n},\bm{y}_{n}\}$ with $X_{n},\bm{y}_{n}$ being the full observation locations and values. Given these observations, the next placement location $\bm{s}_{n}^{*}$ at placement time $t_{n}$ is decided by
@@
To compute the expectation in (2), one could first approximate the full posterior Gaussian process to the marginal posterior predictive distribution over a sufficiently fine spatio-temporal grid $R\times\mathcal{T}$, which makes the distribution multivariate Gaussian. Subsequently, we find the covariance matrix of the marginal posterior predictive and calculate its log determinant for entropy evaluation.
The objective function of (2), following the approach above, admits a closed-form expression but is costly to compute. Fortunately, using the symmetric property of the mutual information, one can obtain an equivalent formulation of the information gain utility that is computationally cheap and exactly measures the entropy of the full posterior. We have
@@ 4 BALLAST
To faithfully measure the utility of a placed drifter, we propose BALLAST: Bayesian Active Learning with Look-ahead Amendment for Sea-drifter Trajectories, a novel algorithm that adjusts the utility computation via look-aheads using vector fields sampled from posteriors. A preliminary version addressing only stationary vector fields appeared as a NeurIPS workshop paper (Zhang et al., 2024).
@@
To compute the acquisition in (4), we approximate the outer expectation over $F\sim p(f|\mathcal{D}_{n})$ using the Monte Carlo method (Robert and Casella, 1999) by taking $J$ draws $F^{(1)},F^{(2)},\ldots,F^{(J)}$ from the posterior $p(f|\mathcal{D}_{n})$ and computing the integrand individually. The choice of $J$ is a key tuning parameter, which we pick $J=20$ as the default following our ablation study’s result in Sections 5.1 and G.
@@
The main computational challenge of the acquisition function of (5) is the GP posterior sampling of $F^{(1)},\ldots,F^{(J)}$ at the $N_{\text{space}}$ spatial grid and $N_{\text{sampT}}$ temporal grid, thus $N_{\text{samp}}=N_{\text{space}}N_{\text{sampT}}$ test points, which costs $O(N_{\text{samp}}^{3})=O(N_{\text{space}}^{3}N_{\text{sampT}}^{3})$ each sample assuming we have obtained the posterior predictive’s covariance matrix. This challenge is tackled below.
4.1 Efficient Sampling from Spatio-Temporal GP Posteriors using the SPDE Approach
To overcome the computational challenge of BALLAST acquisition, we present a computationally efficient posterior sampling method by leveraging the dynamic formulation of spatio-temporal GPs of the SPDE approach (Sarkka et al., 2013; Solin, 2016).
Under the SPDE formulation, 1D Matérn GP can be cast as the solution of a linear SDE, which allows inference using the Kalman filter and Rauch-Tung-Striebel (RTS) smoother that costs linearly in time (Hartikainen and Särkkä, 2010). When the GP has a separable spatial kernel in addition to the Matérn temporal kernel, the same tools can be applied (Sarkka et al., 2013) and cost linearly in time and cubically in space. See Section F for more details.
Our temporal Helmholtz model belongs to the class of GP models admitting such SPDE-GP formulation. However, such a formulation almost always works with gridded observations with overlapping observations and test spatial grids. This is not the case here, as Lagrangian observations are non-gridded. Direct implementation would thus yield prohibitive extra computational cost, making any potential speed-up futile. Details of this cost analysis can be found in F.4.
Below, we present a method to achieve this by adjusting the kernel that enables regression via standard GP and posterior prediction and sampling via the SPDE approach without additional approximations.
Our proposed method considers the extended GP $\bm{f}=[f,\partial_{t}f]^{T}$ with $f\sim\mathcal{GP}(0,k_{\text{tHelm}})$ and regresses the observations with it. In particular, using the properties of kernels under linear operators (Agrell, 2019), the extended GP would have the following kernel
@@ 4.1 Efficient Sampling from Spatio-Temporal GP Posteriors using the SPDE Approach
Since $k_{\text{tHelm}}$ is separable, the partial derivatives are only w.r.t. the Matèrn temporal kernels [^1] To implement such a GP using automatic differentiation, one may need to manually define the Matérn kernel to avoid wrong derivative values w.r.t the kernel’s distance function as most common Python GP packages such as GPJax (Pinder and Dodd, 2022) clip the distance function near zero for numerical stabilities..
To sample vector fields for BALLAST at time $t_{n}$ with observations $\mathcal{D}_{n}$, we first use the extended posterior distribution $\bm{f}|\mathcal{D}_{n}$ to generate the SPDE initial condition by drawing from $\bm{f}(R,t_{n})|\mathcal{D}_{n}$, then propagate the initial condition until terminal time $T$ using the state space model. This approach maintains the linear in time complexity of the SPDE sampling while avoiding the need to filter over the extended spatial locations.
@@ 4.2 BALLAST Algorithm
where the blue indicates the complexity that can be reduced using parallelization. It is noteworthy that the cost of one iteration of BALLAST-EIG without the technique introduced in Section 4.1 is $O(JN_{\text{space}}^{3}N_{\text{sampT}}^{3})$ – a far greater cost considering both $N_{\text{space}}$ and $N_{\text{sampT}}$ are large for practical deployments.
Our proposed Algorithm 1 builds on a separable, spatio-temporal GP surrogate with the temporal kernel being Matérn for the execution of the SPDE sampling procedure described in Section 4.1. This could be weakened following the work of Solin (2016) on constructing SPDE formulations for a broader range of kernels. However, the core BALLAST mechanism of trajectory projection is compatible with any utility choice and active learning surrogate models.
@@
1:Deployment number $M$. BALLAST sample number $J$. Temporal Helmholtz GP $f$. Spatial grid $R$. Terminal time $T$.
@@
4:  Optimise GP hyperparameters using existing observations.
5:  for $j=1,2,\ldots,J$ do $\triangleright$ parallelizable
@@
9:   for $\bm{s}\in R$ do $\triangleright$ parallelizable
10:     Simulate the trajectory of a newly placed observer at $\bm{s}$.
@@
13:  Aggregate the utility contributions from the $J$ sampled vector fields to obtain the next placement location $\bm{s}^{*}_{m}$ using (5).
@@ 5 Experiments
After an ablation study empirically analysing the tuning parameter choice of the BALLAST policy, we investigate the effectiveness of BALLAST for Lagrangian observer placement under ground truth generated by the temporal Helmholtz GP surrogate and under the high-fidelity Stanford Unstructured Nonhydrostatic Terrain-following Adaptive Navier–Stokes Simulator (SUNTANS, Fringer et al. (2006)) numerical model.
@@
The deployment policy proposed by Chen et al. (2024b) works under the Lagrangian data assimilation inference framework, and considers two criteria: (1) “the drifters are deployed at locations where they can travel long distances”, and (2) “place the drifters at locations that are separate from each other”. As we are working under a different inference framework, we adapt their policy and compute the criteria using GP posteriors and BALLAST samples, which gives us the DIST-SEP policy. Implementation details of all considered polices can be found in Section H.3.

### v2

@@ Abstract
We introduce a formal active learning methodology for guiding the placement of Lagrangian observers to infer time-dependent vector fields – a key task in oceanography, marine science, and ocean engineering – using a physics-informed spatio-temporal Gaussian process surrogate model. The majority of existing placement campaigns either follow standard ‘space-filling’ designs or relatively ad-hoc expert opinions. A key challenge to applying principled active learning in this setting is that Lagrangian observers are continuously advected through the vector field, so they make measurements at different locations and times. It is, therefore, important to consider the likely future trajectories of placed observers to account for the utility of candidate placement locations. To this end, we present BALLAST: Bayesian Active Learning with Look-ahead Amendment for Sea-drifter Trajectories. We observe noticeable benefits of BALLAST-aided sequential observer placement strategies on both synthetic and high-fidelity ocean current models. In addition, we developed a novel GP inference method – the Vanilla SPDE Exchange (VaSE) – to boost the GP posterior sampling efficiency, which is also of independent interest.
@@ 1 Introduction
After realising the inadequacy of standard active learning methods for Lagrangian observers (see Section 3.1 for more details), we propose BALLAST — Bayesian Active Learning with Lookahead Amendment for Sea-drifter Trajectories. BALLAST accounts for the data structure of Lagrangian observers by simulating hypothetical trajectories using vector fields. Here, we used a spatio-temporal vector-output Gaussian process (GP) surrogate (see Figure 1 for an illustration) and proposed an information-theoretic utility for the active learning, and devised an original GP inference method, the Vanilla SPDE Exchange (VaSE), for efficient posterior sampling that is thousands of times faster than SPDE-GP and billions of times faster than standard GP for the problems we consider here (see Section 4.1 for details).
Our contributions can be summarised as follows: (i) we introduce active learning concepts to the literature of Lagrangian observer placements, (ii) we propose BALLAST, a novel active learning amendment that accounts for Lagrangian observations using samples from surrogates, and (iii) we develop the vanilla-SPDE exchange, a new GP inference method combining standard GP regression and the SPDE approach (Sarkka et al., 2013), for efficient BALLAST utility computation, which may be of independent interest especially for spatio-temporal GPs with non-gridded observations. Our numerical results suggest noticeable benefits of BALLAST-aided active learning for sequential observer deployment on both synthetic and high-fidelity ocean current models.
@@ 2 Background
2.1 Gaussian Process for Time-Dependent Vector Fields
A surrogate model is employed in active learning to model the unknown object of interest $f$ while providing uncertainty quantification, commonly chosen to be a GP (Williams and Rasmussen, 2006; Gramacy, 2020). In this study the system of interest is a time-dependent vector field, so we consider a spatio-temporal, vector-output GP as the surrogate used for active learning.
Following Ponte et al. (2024), we consider an extension to the Helmholtz kernel $k_{\text{Helm}}$ of Berlinghieri et al. (2023) – a vector-output kernel (Alvarez et al., 2012) using the Helmholtz decomposition (Bhatia et al., 2012) to more realistically portray the vector field structure – by including a separable temporal kernel $k_{\text{time}}$, which results in the temporal Helmholtz kernel $k_{\text{tHelm}}\left((\bm{s},t),(\bm{s}^{\prime},t^{\prime})\right)=k_{\text{Helm}}(\bm{s},\bm{s}^{\prime})k_{\text{time}}(t,t^{\prime})$ with $\bm{s},\bm{s}^{\prime}\in\operatorname{\mathbb{R}}^{2},t,t^{\prime}\in\operatorname{\mathbb{R}}$. The spatial Helmholtz kernel $k_{\text{Helm}}$ is constructed from linear differential operators applied to the potential and stream function kernels, which are two two-dimensional input, scalar output kernels such as the SE kernel (see Section B.2 for details). The temporal kernel $k_{\text{time}}$ is set to be a Matérn $3/2$ kernel: evidence in oceanographic research suggests that the smoothness $\nu$ is around 2; as this leads to a non-analytic representation of the Bessel function, often $\nu=3/2$ is taken as a sufficient approximation (Lilly et al., 2017; Ponte et al., 2024). We will also use $\bm{x}=(\bm{s},t)$ to denote the input of the GP. An illustration of the spatio-temporal GP regression of Lagrangian trajectories is displayed as Figure 1.
@@ 2.2 Sequential Experimental Design
Sequential experimental design, with active learning as a special case, selects an optimal measurement point $x^{*}_{n}$ from measurement set $X$ at each time $t_{n}$ using existing data $\mathcal{D}_{n}$ by optimising the expectation of utility function $U$ over the posterior predictive distribution $p(y|\mathcal{D}_{n},x)$ at $x\in X$, mathematically formulated as
@@
2.3 Problem Setup
Here, our active learning object of interest $f$ depends on both space and time. We assume pre-determined decision times, so we only select the placement location of the Lagrangian observers at each decision time. This simplifying assumption is commonly imposed in the literature to reduce the sequential design’s search space (Huang et al., 2022; Fanshawe and Diggle, 2012; Wikle and Royle, 1999), and is an accurate representation of many real-world drifter deployment practices (Lilly and Pérez-Brunius, 2021). Additionally, we assume the Lagrangian observers produce data with negligible delay – a realistic assumption as modern surface drifters broadcast GPS positions in real time at regular intervals with high accuracy (Novelli et al., 2017).
@@ 3 Active Learning of Time-Dependent Vector Fields
and $\sigma_{\text{obs}}$ is the standard deviation of the observation noise. The active learning selects the placement location of Lagrangian observers at each placement time in order to maximise the information gain of the surrogate $f$ over the full spatio-temporal grid.
@@
Let $t_{n}$ be the time when we are deciding the placement location of the $n$-th observer. The observations so far are denoted by $\mathcal{D}_{n}=\{X_{n},\bm{y}_{n}\}$ with $X_{n},\bm{y}_{n}$ being the full observation locations and values. Given these observations, the next placement location $\bm{s}_{n}^{*}$ at placement time $t_{n}$ is
@@
Although standard computation of expected information gain like (2) involves a computationally prohibitive cost (Ryan et al., 2016), we can compute our expected utility efficiently here using the properties of Gaussian distributions’ entropies as well as the symmetry of mutual information.
In particular, the posterior predictive distribution of a GP over a spatio-temporal grid, i.e. $H(p(f(R\times\mathcal{T})|\mathcal{D}))$, is a multivariate Gaussian and its entropy is directly computable analytically from the distribution’s covariance matrix. Furthermore, this entropy has no dependency on the observation value, thus the expected entropy is identical to the entropy itself. Finally, using the symmetry of mutual information, one can obtain an equivalent formulation of the information gain in (2) that is computationally cheaper than its original form. We remark that such a reformulation using mutual information symmetry has been applied to entropy search (Hennig and Schuler, 2012) to yield the more cost-efficient predictive entropy search (Hernández-Lobato et al., 2014).
Overall, we have
@@ 4 BALLAST
To faithfully measure the utility of a placed drifter, we propose BALLAST: Bayesian Active Learning with Look-ahead Amendment for Sea-drifter Trajectories, a novel algorithm that adjusts the utility computation via look-aheads using vector fields sampled from posteriors.
@@
To compute the acquisition in (4), we approximate the outer expectation over $F\sim p(f|\mathcal{D}_{n})$ using the Monte Carlo method (Robert and Casella, 1999) by taking $J$ draws $F^{(1)},F^{(2)},\ldots,F^{(J)}$ from the posterior $p(f|\mathcal{D}_{n})$ and computing the integrand individually. The choice of $J$ is a key tuning parameter, which we pick $J=20$ as the default following our ablation study’s result in Sections 5.1 and H.
@@
Close inspections of (5) indicate the computational bottleneck of the optimisation is the posterior sampling of vector fields $F^{(1)},\ldots,F^{(J)}$ over the full spatio-temporal grid. Standard GP posterior sampling scales cubically in test points as it involves computing the matrix square root of a Gram matrix. Here, our test points consist of the full spatio-temporal grid of size $N_{\text{samp}}=N_{\text{space}}N_{\text{sampT}}$ – with $N_{\text{space}}$ spatial points and $N_{\text{sampT}}$ temporal points – commonly of the magnitude $10^{5}$ or higher, making standard sampling computationally impossible. In Section 4.1 below, we propose a novel GP inference method to overcome this bottleneck.
4.1 Vanilla-SPDE Exchange
To solve the challenge of posterior sampling for BALLAST, we propose the Vanilla SPDE Exchange (VaSE), a novel GP inference method that synergises the computational benefits of the standard GP and SPDE (Solin, 2016) frameworks, which is also of independent interest.
Under the SPDE-GP re-formulation, a 1D Matérn GP can be cast as the stationary solution of a linear SDE, which allows inference using the Kalman filter and Rauch-Tung-Striebel (RTS) smoother that costs linearly in time (Hartikainen and Särkkä, 2010). When the GP has a separable spatial kernel in addition to the Matérn temporal kernel, the same tools can be applied (Sarkka et al., 2013) and cost linearly in time and cubically in space. See Section F for more details.
While the SPDE approach of Sarkka et al. (2013) offers a computationally efficient re-formulation for spatio-temporal GPs, the method is not suitable for situations where the observation locations and test locations are minimally overlapping. Since the spatial component of the GP is included in the SPDE using the Gram matrix of all considered locations, near-distinct observation and test locations yield a huge Gram matrix and cost cubically in size. As our Lagrangian observations are non-gridded, predicting on gridded test locations would thus yield prohibitive extra computational cost, making any potential speed-up of SPDE-GP futile. Details of this cost analysis can be found in F.4.
VaSE, on the other hand, uses standard GP regression to bypass the need to consider observation locations for SPDE-GP, so as to achieve efficient posterior sampling. We consider the extended GP $\bm{f}=[f,\partial_{t}f]^{T}$ with $f\sim\mathcal{GP}(0,k_{\text{tHelm}})$ and regress the observations with it. In particular, using the properties of kernels under linear operators (Agrell, 2019), the extended GP has kernel
@@ 4.1 Vanilla-SPDE Exchange
Since $k_{\text{tHelm}}$ is separable, the partial derivatives are only w.r.t. the Matérn temporal kernels.
To sample vector fields for BALLAST at time $t_{n}$ with observations $\mathcal{D}_{n}$ using VaSE, we first use the extended posterior distribution $\bm{f}|\mathcal{D}_{n}$ to generate the SPDE initial condition by drawing from $\bm{f}(R,t_{n})|\mathcal{D}_{n}$, then propagate the initial condition until terminal time $T$ using the state space model. Full details of VaSE can be found in Section G.
Method Cost
Standard $O(N_{\text{obs}}^{3}+N_{\text{pred,s}}^{3}N_{\text{pred,t}}^{3}+N_{\text{obs}}^{2}N_{\text{pred,s}}N_{\text{pred,t}}+N_{\text{obs}}N_{\text{pred,s}}^{2}N_{\text{pred,t}}^{2}))$
SPDE $O((N_{\text{obs}}+N_{\text{pred,s}})^{3}N_{\text{obs,t}}+N_{\text{pred,s}}^{2}N_{\text{pred,t}})$
VaSE $O(N_{\text{obs}}^{3}+N_{\text{pred,s}}^{2}N_{\text{obs}}+N_{\text{pred,s}}N_{\text{obs}}^{2}+N_{\text{pred,s}}^{2}N_{\text{pred,t}})$
Table 1: Computational cost comparison of standard, SPDE, VaSE for drawing one posterior sample. We highlight the dominating terms $N_{\text{pred,s}}^{3}N_{\text{pred,t}}^{3}$ for standard and $(N_{\text{obs}}+N_{\text{pred,s}})^{3}N_{\text{obs,t}}$ for SPDE that make the two traditional methods drastically more expensive than VaSE. Full details in Section G.1.
The computational costs of drawing one posterior sample using the three approaches (standard GP, SPDE-GP, vanilla-SPDE exchange) are compared in Table 1, where $N_{\text{obs}}$ ($\approx 200$) denotes the observation number, $N_{\text{obs,t}}$ ($\approx 200$) denotes the number of distinct observation times, $N_{\text{pred,s}}$ ($\approx 500$) and $N_{\text{pred,t}}$ ($\approx 1000$) denotes the number of prediction spatial and temporal points. Using the approximate values, we can notice that VaSE is of the cost order $10^{8}$, significantly cheaper than the $10^{11}$ cost of SPDE and $10^{17}$ cost of standard GP.
@@ 4.2 BALLAST Algorithm
where the blue indicates the complexity that can be reduced using parallelization.
Our proposed Algorithm 1 builds on a separable, spatio-temporal GP surrogate with the temporal kernel being Matérn for the execution of the SPDE sampling procedure described in Section 4.1. This could be weakened following the work of Solin (2016) on constructing SPDE formulations for a broader range of kernels. However, the core BALLAST mechanism of trajectory projection is compatible with any utility choice and active learning surrogate models. We also remark that we initialise the algorithm with a uniformly drawn location for generality and simplicity.
@@
1:Input: Deployment number $M$. BALLAST sample number $J$. Temporal Helmholtz GP $f$. Spatial grid $R$. Terminal time $T$.
@@
4:  Optimise GP hyperparameters using observations.
5:  for $j=1,2,\ldots,J$ (parallelizable) do
@@
9:   for $\bm{s}\in R$ (parallelizable) do
10:    Simulate the trajectory starting at $\bm{s}$.
@@
13:  Aggregate the utility contributions from $J$ samples to obtain the next placement $\bm{s}^{*}_{m}$ via (5).
@@ 5 Experiments
After an ablation study empirically analysing the tuning parameter choice of the BALLAST policy, we investigate the effectiveness of BALLAST for Lagrangian observer placement under ground truth generated by the temporal Helmholtz GP surrogate and under the high-fidelity Stanford Unstructured Nonhydrostatic Terrain-following Adaptive Navier–Stokes Simulator (SUNTANS) Fringer et al. (2006).
@@
The deployment policy proposed by Chen et al. (2024b) works under the Lagrangian data assimilation inference framework, and considers two criteria: (1) “the drifters are deployed at locations where they can travel long distances”, and (2) “place the drifters at locations that are separate from each other”. As we are working under a different inference framework, we adapt their policy and compute the criteria using GP posteriors and BALLAST samples, which gives us the DIST-SEP policy. Implementation details of all considered polices can be found in Section I.3, and the anonymized Github repo for the codes used in the experiments can be found at https://github.com/ShuSheng3927/BALLAST.